OpenStudy (anonymous):

Determine whther the series is convergent or divergent. I solved it, But I got a different answer.

OpenStudy (anonymous):

\[\sum_{i=1}^{\infty}=\frac{ \ln n }{n ^{3} }\]

OpenStudy (anonymous):

that would converge

OpenStudy (anonymous):

at a certain point, the denominator would be so large, that it would almost be like adding zero

OpenStudy (anonymous):

Nope, this is divergent, because your index is i, not n :))

OpenStudy (anonymous):

But in all seriousness, if you let me fix it for you... \[\Large \sum_{n=1}^\infty\frac{\ln(n)}{n^3}\]

OpenStudy (anonymous):

@PeterPan Thanks... I know that it is convergent... WHat value did you get?

OpenStudy (anonymous):

What value? I didn't know you were supposed to evaluate the sum itself... that's so hard D:

OpenStudy (anonymous):

well we have to... I got a different answer from the text book... I just wan to see if I got the value right...

OpenStudy (anonymous):

How did you get a sum, then?

OpenStudy (anonymous):

integrate then find the limit

OpenStudy (anonymous):

Integrate?! But I thought this was a sum? You mean like this? \[\Large \int\limits_1^\infty \frac{\ln(x)}{x^3}dx\]

OpenStudy (anonymous):

yes...but the initial value of x must be 2.

OpenStudy (anonymous):

Why 2?

OpenStudy (anonymous):

It starts with 1, doesn't it?

OpenStudy (anonymous):

because that is where the series is decreasing

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

But you can very well integrate it from 1, can't you? \[\Large = \int\limits_0^\infty ue^{-2u}du\]

OpenStudy (anonymous):

\[\Large =\left. -\frac{e^{-2x}(2x+1)}4\right]_0^\infty\]

OpenStudy (anonymous):

Which is \(\Large \frac34\) I believe...

OpenStudy (anonymous):

why did you start at zero?

OpenStudy (anonymous):

Oh, because I kind of let u = ln(x) So, you start u from ln(1) which is zero.

OpenStudy (anonymous):

ok...thanks.. it is 1/4

OpenStudy (anonymous):

1/4 is the answer in the book?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oops... yeah, it is 1/4 my mistake, 2(0) + 1 is most definitely not 3/4 See, this is why people shouldn't count on me ^.^

OpenStudy (anonymous):

aww.... :< Because you are Peter pan...

OpenStudy (anonymous):

I make these annoying arithmetic error after an awesome calculus demonstration... still... I AM awesome >:)

OpenStudy (anonymous):

\[\Large =\left. -\frac{e^{-2x}(2x+1)}4\right]_0^\infty= \frac{e^{-2(0)}[2(0)+1]}{4}=\frac14\]

OpenStudy (anonymous):

My actual name. Pan (Korean: 판) is my middle name ^.^

OpenStudy (anonymous):

A quarter Korean, the rest is British.

OpenStudy (anonymous):

nice...pm?

OpenStudy (anonymous):

pm?

OpenStudy (anonymous):

private message b.c. this is personal... Just learned that shortcut when I was here